Evaluate an investment projects and choose best investment option.
Bank deposit rate 6% per annum:
1) Investment 2000, gives a profit at the end of the 1st year 200, 2nd year.b. 500, 3.g.b. EUR 1200;
2) a bon (4 coupons)  with 500  EUR face value. The annual benefit (coupon payment) is 50 EUR. The term of the  bond is 4 years.
1
Expert's answer
2021-01-11T12:02:38-0500

If bank deposit rate 6% per annum, then:

1)

"NPV = -2,000 + 200\/1.06 + 500\/1.06^2 + 1,200\/1.06^3 = -358.78."

2) "NPV = -2,000 + 200\/1.06 + 200\/1.06^2 + 200\/1.06^3 + 200\/1.06^4 = -1,306.98."

So, the first option is better.